OpenStudy (anonymous):

I'll put the logarithm bases into () . Solve log(2)log(4)log(8)64 .

OpenStudy (anonymous):

64 = 8^2 so log(8)64 = 2 log(2)log(4)2 2 = 4^0.5 so log(4)2 = 0.5 log(2)0.5 0.5 = 2^-1 so log(2)0.5 = -1

OpenStudy (anonymous):

Thanks !